OpenStudy (czarluc):

sinx+sinx=0? is it 2sinx? or sin2x?

OpenStudy (solomonzelman):

sinx+sinx=0 sinx=-sinx sin(x)=sin(-x) x=-x

OpenStudy (solomonzelman):

x=-x there is only one value that makes it true.

OpenStudy (solomonzelman):

2sinx=sinx+sinx if you mean that...

OpenStudy (anonymous):

2sin x=0 implies sin x =0 implies x=integral multiple of pie

OpenStudy (czarluc):

when will it equal to sin2x?

OpenStudy (czarluc):

in what kind of situation?

OpenStudy (anonymous):

sin 2 xis 2sinx cosx

OpenStudy (solomonzelman):

no, 2sinx sin(2x) is different sin(2x)=sin(x)cos(x)+sin(x)cos(x)=2sin(x)cos(x)

OpenStudy (czarluc):

oh ok , uhm another question... what are the solutions of this? 4cos2x=8sinxcosx

OpenStudy (solomonzelman):

4cos(2x)=8sinxcosx cos(2x)=2sinxcosx cos(2x)=sin(2x)

OpenStudy (anonymous):

integral multiple of pie is the solution that is npie, where n is an integer

OpenStudy (solomonzelman):

2x=45 x=?

OpenStudy (czarluc):

where did you get 45?

OpenStudy (czarluc):

thanks princer and solomon

OpenStudy (solomonzelman):

 45 degrees is the only time when cosine and sine are equal.

OpenStudy (solomonzelman):

so it is 45/2

OpenStudy (solomonzelman):

the x is 45/2 I mean

OpenStudy (czarluc):

what would happen in the sin2x=cos2x?

OpenStudy (solomonzelman):

you will need to tell me first the angle at which sin(a)=cos(a) where a=2x

OpenStudy (solomonzelman):

I think I have disconnected about 20 times for the past 10-15 minutes.

OpenStudy (czarluc):

it's not indicated?

OpenStudy (czarluc):

it's fine lol

OpenStudy (solomonzelman):

the angle at which cos=sin is 45, so to make 2x =45, the x has to be 45/2.

OpenStudy (czarluc):

so what would be the values?

OpenStudy (solomonzelman):

your interval is (-infinity,+infinity) or [0,360] or some other interval?

OpenStudy (czarluc):

[0,360] is the interval

OpenStudy (solomonzelman):

so, your solution will be pi/4

OpenStudy (czarluc):

only pi/4?

OpenStudy (solomonzelman):

or 135º

OpenStudy (czarluc):

why 135? aren't there two functions?

OpenStudy (solomonzelman):

my bad, 135 is not a solution, because cos(135) doesn't equal sin(135).

OpenStudy (solomonzelman):

x should be 45.

OpenStudy (solomonzelman):

no other coterminal angles, because the interval is [0,360] (only one round)

OpenStudy (solomonzelman):

I mean 2x is 45

OpenStudy (czarluc):

oh ok so entering the values unto sin/cos is for checking?

OpenStudy (solomonzelman):

I think I should get a rest.. I am sorry.

OpenStudy (czarluc):

it's fine. thanks btw